40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 17 01:55:16 2015, ericq, Update, LSC, Uncooperative AUX X 
    Reply  Fri Dec 18 17:37:04 2015, rana, Update, LSC, Uncooperative AUX X 
       Reply  Mon Dec 21 02:29:49 2015, ericq, Update, LSC, AUX X RIN measurements auxXRIN.pdf
          Reply  Tue Jan 5 02:54:38 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
             Reply  Tue Jan 5 13:17:06 2016, rana, Update, LSC, AUX X Freq Noise attempt 
                Reply  Tue Jan 5 16:33:45 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
                   Reply  Thu Jan 7 04:28:39 2016, ericq, Update, LSC, AUX X Freq Noise measured PLLspec.pdfPLL_OLG.pdf
                      Reply  Thu Jan 7 16:52:32 2016, ericq, Update, LSC, AUX X Freq Noise measured AUXfreqnoise.pdfauxXmeasurements.zip
                         Reply  Thu Jan 7 19:04:25 2016, Koji, Update, LSC, AUX X Freq Noise measured 
                            Reply  Fri Jan 8 14:47:33 2016, ericq, Update, LSC, AUX Y Freq Noise measured AUX_freqnoises.pdfAUXY_Jan8.zip
                               Reply  Fri Jan 8 20:02:49 2016, rana, Update, LSC, AUX Y Freq Noise measured 
                               Reply  Sat Jan 9 00:39:15 2016, gautam, Update, LSC, AUX Y Freq Noise re-measured AuxPLL.pdf
                                  Reply  Mon Jan 11 19:01:56 2016, gautam, Update, LSC, PLL Marconi Investigation 6x
                                     Reply  Tue Jan 12 03:03:55 2016, ericq, Update, LSC, Frequently making noise 
                                        Reply  Tue Jan 12 19:38:31 2016, gautam, Update, LSC, Frequently making noise 2016_01_AUXLaser.tar.gzOLGs.pdfvariedCurrent.pdfvariedTemp.pdfPLL_setup.pdf
Message ID: 11889     Entry time: Thu Dec 17 01:55:16 2015     Reply to this: 11892
Author: ericq 
Type: Update 
Category: LSC 
Subject: Uncooperative AUX X 

[ericq, Gautam]

We were not able to fix the excess frequency noise of the AUX X laser by the usual laser diode current song and dance. Unfortunately, this level of noise is much too high to have any realistic chance of locking.  angry

We're leaving things back in the IR beat -> phase tracker state with free running AUX lasers, on the off chance that there may be anything interesting to see in the overnight data. This may be limited by our lack of automatic beatnote frequency control. (Gautam will soon implement this via digital frequency counter). I've upped the FINE_PHASE_OUT_HZ_DQ frame rate to 16k from 2k, so we can see more of the spectrum.

For the Y beat, there is the additional weird phenomenon that the beat amplitude slowly oscillates to zero over ~10 minutes, and then back up to its maximum. This makes it hard for the phase tracker servo to stay stable... I don't have a good explanation for this. 

ELOG V3.1.3-